# OneVision: An End-to-End Generative Framework for Multi-view E-commerce Vision Search

### v3

@@ 1. Introduction
Recent years have witnessed substantial progress in vision search, propelled by advances across multiple fronts, including object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and visual representation learning (Lee et al., 2022a; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These developments have driven the construction of large-scale industrial systems, with major e-commerce platforms such as eBay (Yang et al., 2017) and Alibaba (Zhang et al., 2018; Chen et al., 2024a) deploying comprehensive vision search pipelines to support production-scale applications. Despite these advances, traditional pipelines in vision search still face structural limitations: (i) Difficulty in handling multi-view discrepancy. Query images and product images often differ significantly in viewpoint, making it hard for conventional retrieval pipelines to align their visual representations and recall the correct products. (ii) Stage-wise objective misalignment. The optimization goals across three stages collide, making it difficult to reach Pareto optimality in both user experience and conversion. (iii) Heavy multi-stage overhead. The multi-stage architecture introduces substantial computational cost and cross-module communication, increasing latency and complicating system maintenance.
@@ 2.1. Vision Search System
Modern image-based search systems have advanced considerably with progress across core computer vision tasks such as object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and representation learning (Lee et al., 2022a; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These advances have substantially improved feature extraction, enabling the generation of compact and discriminative visual embeddings from regions of interest. Building on such features, large-scale approximate nearest neighbor (ANN) retrieval is used to construct candidate pools, which are subsequently refined and personalized through pre-ranking and ranking (Dagan et al., 2021; Nadkarni and Dasararaju, 2022; Dagan et al., 2023; You et al., 2024).
@@ 2.2. Generative Retrieval System
Encode method in Generative Retrieval. Encoding is a core component of generative retrieval. RQ-VAE(Lee et al., 2022b) encodes images into residual-quantized codes, achieving high capacity with compact codebooks and modeling long-range dependencies efficiently via an autoregressive RQ-Transformer. QARM (RQ-KMeans)(Luo et al., 2024) aligns multimodal representations and compresses them into learnable SIDs through residual K-means, supporting end-to-end training and consistent semantics at scale. FSQ (Mentzer et al., 2023) replaces vector quantization with bounded scalar quantizers. This approach maintains performance while enabling large, well-utilized codebooks and easier optimization, which is particularly useful when reconstruction quality and sampling efficiency are critical.
@@ 3.1. Vision-aligned Residual Quantization
Semantic encoding of images is a key step for generative vision search. A common approach is hierarchical semantic encoding from coarse to fine levels. However, effective SIDs must satisfy two requirements: consistency, where similar items or multi-view images of the same object share shallow-level codes; and uniqueness, where different images remain distinguishable at deeper levels. However, current quantization methods face three major challenges: (i) Techniques like FSQ (Mentzer et al., 2023) and OPQ (Ge et al., 2013) struggle to hierarchically capture core attributes among similar items, limiting generative training; (ii) General methods such as VQ-VAE (Van Den Oord et al., 2017), RQ-VAE (Lee et al., 2022b), and RQ-KMeans (Luo et al., 2024) use shared codebooks, limiting their ability to preserve uniqueness; (iii) Offline image encoding for retrieval often depends on object detection for item encoding, but incorrect main-subject recognition can yield erroneous embeddings and introduce noise.
@@ 3.1.3. VRQ Hierarchical Quantization Encoding
and the residual is updated as $r_{l+1}=r_{l}-e_{c_{l}}^{(l)}$. Finally, we can obtain the SID ($c_{0},...,c_{L-1}$) and the quantized representation at level $l$ is $\hat{f}^{(l)}=\sum_{i=0}^{l}e_{c_{i}}^{(i)}.$ The commitment loss in RQ-VAE (Lee et al., 2022b) is
@@ 3.3. Dynamic Pruning
Inspired by (Omri et al., 2025; Wang et al., 2025b), we incorporate a visual token selection algorithm into the encoder output. Given a query image with $V_{max}$ visual tokens, our method selects a subset $V_{sub}$ of the most relevant tokens for subsequent decoding. Specifically, during both training and inference, we apply K-means++ (Arthur and Vassilvitskii, 2006) clustering on the embeddings of the original $V_{max}$ visual tokens to form $V_{sub}$ cluster centers. Within each cluster, only the token closest to the center is retained. This procedure effectively groups similar patterns, highlights salient regions in the image, and filters out less informative signals. We then adopt a distillation framework, where a frozen post-trained generative model serves as the teacher and the pruned model as the student. KL divergence is used to ensure consistency between their output distributions:
@@ 4.3. More Offline Evaluation
Finally, Figure 5 shows that with curriculum-based visual token pruning, the model can dynamically reduce the number of visual tokens while preserving retrieval accuracy. HR@4 and MRR@4 remain stable even when compressing from the full 197 tokens to only 33 tokens. This compression yields substantial efficiency gains: real-time inference latency (RT) drops by about 20% and memory usage decreases by roughly 13%. By reducing redundancy in the encoder’s visual representations and easing the burden of beam search decoding, this approach markedly lowers online inference cost while maintaining strong semantic understanding.
@@ 4.4. Online A/B Testing
where $\omega_{1},\omega_{2},\omega_{3}$ are tunable weights that control the relative importance of recent clicks, GMV, and completed orders within a 30-day window. This prioritizes items with stronger recent engagement and transaction value within a code cluster. Results show that multi-stage supervised training together with VRQ encoding yields consistent improvements over OnlineMCA: CTR +2.15%, CVR +2.27%, OPM +3.60%, Click +1.80%, and Order +3.12%. In contrast, OnlineMCA w/o ranking exhibits marked declines (e.g., Order -12.65%), underscoring the centrality of ranking in Photo Search and indirectly confirming that OneVision, augmented by its dynamic item sequence, already demonstrates strong ranking capability. Overall, these outcomes indicate that OneVision matches or surpasses the baseline OnlineMCA while simplifying the online stack without inducing seesaw effects, and it shows clear headroom for continued iteration in the era of large generative models.
@@ 5. Further Analysis
We further validate these two points through an in-depth analysis of the online A/B test results. With multi-view modeling, the overall Item CTR shows a clear +1.61% relative gain across all categories. Notably, women’s clothing improves by about +3.4%, while bags & accessories increase by approximately +4.8%. Meanwhile, unified similarity definition and improved category alignment lead to consistent PV CTR gains across category tiers: the top, middle, and long-tail segments increase by +2.01%, +0.82%, and +1.34%, respectively. These results (see Figure 6 and Table 6) highlight OneVision’s effectiveness in strengthening both identical-item retrieval and category-level consistency, ultimately driving measurable online CTR improvements and enhancing user experience.
@@
Method Item CTR PV CTR CVR OPM Order OnlineMCA w/o ranking -5.01% -5.51% -10.73% -10.39% -12.65% OneVision +2.15% +1.80% +2.27% +3.60% +3.12%
@@
Method Top Middle Long-tail OneVision +2.01% +0.82% +1.34%

### v4

@@ 1. Introduction
Recent years have witnessed substantial progress in vision search, propelled by advances across multiple fronts, including object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and visual representation learning (Lee et al., 2022b; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These developments have driven the construction of large-scale industrial systems, with major e-commerce platforms such as eBay (Yang et al., 2017) and Alibaba (Zhang et al., 2018; Chen et al., 2024a) deploying comprehensive vision search pipelines to support production-scale applications. Despite these advances, traditional pipelines in vision search still face structural limitations: (i) Difficulty in handling multi-view discrepancy. Query images and product images often differ significantly in viewpoint, making it hard for conventional retrieval pipelines to align their visual representations and recall the correct products. (ii) Stage-wise objective misalignment. The optimization goals across three stages collide, making it difficult to reach Pareto optimality in both user experience and conversion. (iii) Heavy multi-stage overhead. The multi-stage architecture introduces substantial computational cost and cross-module communication, increasing latency and complicating system maintenance.
@@ 2.1. Vision Search System
Modern image-based search systems have advanced considerably with progress across core computer vision tasks such as object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and representation learning (Lee et al., 2022b; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These advances have substantially improved feature extraction, enabling the generation of compact and discriminative visual embeddings from regions of interest. Building on such features, large-scale approximate nearest neighbor (ANN) retrieval is used to construct candidate pools, which are subsequently refined and personalized through pre-ranking and ranking (Dagan et al., 2021; Nadkarni and Dasararaju, 2022; Dagan et al., 2023; You et al., 2024).
@@ 2.2. Generative Retrieval System
Encode method in Generative Retrieval. Encoding is a core component of generative retrieval. RQ-VAE(Lee et al., 2022a) encodes images into residual-quantized codes, achieving high capacity with compact codebooks and modeling long-range dependencies efficiently via an autoregressive RQ-Transformer. QARM (RQ-KMeans)(Luo et al., 2024) aligns multimodal representations and compresses them into learnable SIDs through residual K-means, supporting end-to-end training and consistent semantics at scale. FSQ (Mentzer et al., 2023) replaces vector quantization with bounded scalar quantizers. This approach maintains performance while enabling large, well-utilized codebooks and easier optimization, which is particularly useful when reconstruction quality and sampling efficiency are critical.
@@ 3.1. Vision-aligned Residual Quantization
Semantic encoding of images is a key step for generative vision search. A common approach is hierarchical semantic encoding from coarse to fine levels. However, effective SIDs must satisfy two requirements: consistency, where similar items or multi-view images of the same object share shallow-level codes; and uniqueness, where different images remain distinguishable at deeper levels. However, current quantization methods face three major challenges: (i) Techniques like FSQ (Mentzer et al., 2023) and OPQ (Ge et al., 2013) struggle to hierarchically capture core attributes among similar items, limiting generative training; (ii) General methods such as VQ-VAE (Van Den Oord et al., 2017), RQ-VAE (Lee et al., 2022a), and RQ-KMeans (Luo et al., 2024) use shared codebooks, limiting their ability to preserve uniqueness; (iii) Offline image encoding for retrieval often depends on object detection for item encoding, but incorrect main-subject recognition can yield erroneous embeddings and introduce noise.
@@ 3.1.3. VRQ Hierarchical Quantization Encoding
and the residual is updated as $r_{l+1}=r_{l}-e_{c_{l}}^{(l)}$. Finally, we can obtain the SID ($c_{0},...,c_{L-1}$) and the quantized representation at level $l$ is $\hat{f}^{(l)}=\sum_{i=0}^{l}e_{c_{i}}^{(i)}.$ The commitment loss in RQ-VAE (Lee et al., 2022a) is
@@ 3.3. Dynamic Pruning
Inspired by (Omri et al., 2025; Wang et al., 2025b), we incorporate a visual token selection algorithm into the encoder output, specifically on the patch-level visual tokens produced by the ViT encoder, before they are fed into the BART encoder. Given a query image with $V_{max}$ visual tokens, our method selects a subset $V_{sub}$ of the most relevant tokens for subsequent decoding. Specifically, during both training and inference, we apply K-means++ (Arthur and Vassilvitskii, 2006) clustering on the embeddings of the original $V_{max}$ visual tokens to form $V_{sub}$ cluster centers. Within each cluster, only the token closest to the center is retained. This procedure effectively groups similar patterns, highlights salient regions in the image, and filters out less informative signals. We then adopt a distillation framework, where a frozen post-trained generative model serves as the teacher and the pruned model as the student. KL divergence is used to ensure consistency between their output distributions:
@@ 4.3. More Offline Evaluation
Finally, Figure 5 shows that with curriculum-based visual token pruning, the model can dynamically reduce the number of visual tokens while preserving retrieval accuracy. HR@4 and MRR@4 remain stable even when compressing from the full 197 tokens to only 33 tokens. This compression yields substantial efficiency gains: real-time inference latency (RT) drops by 17.41% and memory usage decreases by 12.03%. By reducing redundancy in the encoder’s visual representations and easing the burden of beam search decoding, this approach markedly lowers online inference cost while maintaining strong semantic understanding.
@@ 4.4. Online A/B Testing
where $\omega_{1},\omega_{2},\omega_{3}$ are tunable weights that control the relative importance of recent clicks, GMV, and completed orders within a 30-day window. This prioritizes items with stronger recent engagement and transaction value within a code cluster. Results show that multi-stage supervised training together with VRQ encoding yields consistent improvements over OnlineMCA: CTR +3.61%, CVR +2.43%, OPM +4.53%, Click +3.43%, and Order +5.89%. In contrast, OnlineMCA w/o ranking exhibits marked declines (e.g., Order -12.65%), underscoring the centrality of ranking in Photo Search and indirectly confirming that OneVision, augmented by its dynamic item sequence, already demonstrates strong ranking capability. Overall, these outcomes indicate that OneVision matches or surpasses the baseline OnlineMCA while simplifying the online stack without inducing seesaw effects, and it shows clear headroom for continued iteration in the era of large generative models.
@@ 5. Further Analysis
We further validate these two points through an in-depth analysis of the online A/B test results. With multi-view modeling, the overall Item CTR shows a clear +1.27% relative gain across Top 30 categories. Notably, women wear improves by about +3.3%, while bags increase by approximately +4.8%. Meanwhile, unified similarity definition and improved category alignment lead to consistent PV CTR gains across category tiers: the top, middle, and long-tail segments increase by +2.58%, +1.63%, and +15.29%, respectively. These results (see Figure 6 and Table 6) highlight OneVision’s effectiveness in strengthening both identical-item retrieval and category-level consistency, ultimately driving measurable online CTR improvements and enhancing user experience.
@@
Method Item CTR PV CTR CVR OPM Order OnlineMCA w/o ranking -5.01% -5.51% -10.73% -10.39% -12.65% OneVision +3.61% +3.43% +2.43% +4.53% +5.89%
@@
Method Top Middle Long-tail OneVision +2.58% +1.63% +15.29%